# 00footnotetext: *Equal contribution

### v2

@@ Appendix B Metrics
Appendix C Detailed Results
@@
Hawkes-1 56.78 61.81 38.09 90.98 104.06
Hawkes-2 64.04 69.62 30.74 91.26 104.60
Nonstationary Poisson 49.51 49.26 50.56 100.22 99.66
Nonstationary Renewal 42.36 44.41 55.82 96.37 98.48
Self-Correcting 34.61 34.46 66.15 99.40 99.27
Stationary Renewal 71.57 62.86 38.11 106.08 103.38
PUBG 56.40 19.88 19.93 76.43 41.06
Reddit Comments 237.27 8.75 15.15 275.13 24.51
Reddit Posts 956.10 0.10 23.73 1091.55 24.41
Taxi 81.71 7.13 16.37 98.53 23.19
Twitter 11.53 21.48 2.94 15.27 24.05
Yelp Airport 21.80 15.61 8.40 30.84 24.53
Yelp Mississauga 43.79 13.28 10.60 56.99 24.64
Mean 132.88 31.43 28.97 171.47 61.22
Total 193.29 232.68

### v3

@@ Appendix C Metrics
Appendix D Ablations
We ablate the hyperparameters $\delta$, $b_{\text{ins}}$ and $b_{\text{sub}}$ in Figs. 5, 6, 7 and 8.
Figure 5: Mean and standard error of $d_{l}$, $d_{\mathrm{IET}}$ and MMD on two datasets as we vary the $\delta$ parameter for substitutions.
Figure 6: Mean and standard error of $d_{l}$, $d_{\mathrm{IET}}$ and MMD on two datasets as we vary the number of insertion bins $b_{\text{ins}}$.
Figure 7: Mean and standard error of $d_{l}$, $d_{\mathrm{IET}}$ and MMD on two datasets as we vary the number of substitution bins $b_{\text{sub}}$.
Figure 8: Mean and standard error of $d_{l}$, $d_{\mathrm{IET}}$ and MMD on two datasets as we vary the number of insertion and substitution bins together.
Appendix E Detailed Results
E.1 Inpainting vs. Forecasting
To demonstrate the flexibility of EdiTPP for conditional generation, we evaluate its performance when generating events on the interval [T/3, 2T/3][T/3,\,2T/3]. In the forecasting setting, the model is conditioned only on events occurring before $T/3$, whereas in the inpainting setting, it is conditioned on both the past ($t<T/3$) and the future ($t>2T/3$). As the results show, providing both past and future context substantially improves the quality of the generated middle segment compared to conditioning on the past alone.
PG R/C R/P Tx Tw Y/A Y/M
$d_{\text{xiao}}$
Inpainting 2.22 ± 0.04 22.66 ± 0.77 13.13 ± 0.35 3.02 ± 0.20 1.58 ± 0.03 0.59 ± 0.02 1.11 ± 0.04
Forecasting 2.27 ± 0.05 25.53 ± 1.16 18.09 ± 0.81 3.27 ± 0.33 1.65 ± 0.04 0.63 ± 0.03 1.13 ± 0.05
MRE
Inpainting 0.29 ± 0.01 5.79 ± 0.96 0.20 ± 0.01 0.60 ± 0.22 1.96 ± 0.09 0.48 ± 0.01 0.74 ± 0.11
Forecasting 0.30 ± 0.01 5.07 ± 1.13 0.33 ± 0.01 0.81 ± 0.13 2.05 ± 0.11 0.52 ± 0.05 0.78 ± 0.08
$d_{\text{IET}}$
Inpainting 0.39 ± 0.01 0.40 ± 0.01 0.01 ± 0.00 0.10 ± 0.01 1.13 ± 0.02 0.80 ± 0.05 0.70 ± 0.06
Forecasting 0.41 ± 0.01 0.44 ± 0.03 0.02 ± 0.00 0.10 ± 0.00 1.18 ± 0.01 0.78 ± 0.07 0.72 ± 0.08
E.2 Parametric TPP samples
To illustrate how well each model captures parametric TPPs, we draw 200 samples for the Hawkes and Self-Correcting processes. In Fig. 10, we plot the cumulative count $N(t)$ for each sample, while Fig. 9 shows each event sequence as a separate row, directly visualizing the events over time. These visualizations further highlight the strong unconditional sampling performance of EdiTPP demonstrated in Table 1.
Figure 9: Event times for 200 samples from ground truth data (GT) and each model. Each event sequence is represented as a separate row.
Figure 10: $N(t)$ for 200 samples from ground truth data (GT) and each model.
E.3 Full results
@@
H1 55.05 ± 28.6 65.93 ± 10.5 37.74 ± 10.6 92.58 ± 34.3 102.50 ± 10.6
H2 63.69 ± 32.4 71.72 ± 9.6 30.98 ± 8.7 97.57 ± 38.9 101.30 ± 10.2
NSP 49.59 ± 7.3 49.66 ± 7.1 50.56 ± 8.2 100.14 ± 9.7 100.16 ± 9.8
NSR 42.39 ± 5.8 44.23 ± 7.2 55.88 ± 7.8 97.58 ± 7.5 100.51 ± 10.5
PG 56.69 ± 7.2 19.71 ± 4.4 19.96 ± 4.8 76.32 ± 8.7 40.90 ± 6.4
R/C 247.83 ± 251.3 8.69 ± 6.6 15.38 ± 7.4 274.49 ± 254.3 24.45 ± 5.7
R/P 972.89 ± 300.2 0.11 ± 0.3 23.74 ± 5.0 1109.78 ± 307.1 24.56 ± 7.4
SC 34.93 ± 5.7 34.37 ± 6.9 66.18 ± 8.6 98.95 ± 7.8 99.61 ± 10.2
SR 70.67 ± 21.6 64.09 ± 11.3 38.00 ± 11.5 108.85 ± 31.4 101.92 ± 10.9
Tw 11.83 ± 9.4 22.34 ± 4.6 3.04 ± 2.3 14.37 ± 10.3 24.46 ± 5.2
Tx 96.62 ± 14.4 9.35 ± 3.3 17.23 ± 4.8 97.69 ± 17.4 24.37 ± 5.2
Y/A 29.43 ± 6.4 22.64 ± 5.0 9.00 ± 3.4 30.42 ± 6.3 24.30 ± 4.8
Y/M 54.81 ± 13.8 17.14 ± 4.2 11.37 ± 3.8 56.54 ± 15.5 24.45 ± 4.9
Mean 137.42 33.08 29.16 173.48 61.04
Total 199.65 234.52